### v1

# Intuitive Fine-Tuning: Towards Unifying SFT and RLHF into a Single Process

Abstract
Supervised Fine-Tuning (SFT) and Reinforcement Learning from Human Feedback (RLHF) are two fundamental processes for enhancing the capabilities of Language Models (LMs) post pre-training, aligning them better with human preferences. Although SFT advances in training efficiency, RLHF delivers better alignment, thus they are often combined. However, common practices simply apply them sequentially without unifying their optimization targets, resulting in trade-offs between fitting different objectives. This approach ignores the opportunities to bridge the paradigm gap and take the strengths from both. To obtain a unified understanding, we interpret SFT and RLHF with two sub-processes — Preference Estimation and Transition Optimization — defined at token level within the Markov Decision Process (MDP) framework. This modeling shows that SFT is only a specialized case of RLHF with inferior estimation and optimization. RLHF evaluates the quality of model’s entire generated answer, whereas SFT only scores predicted tokens based on preceding tokens from target answers. Therefore, SFT overestimates the ability of model, leading to inferior optimization. Building on this view, we introduce Intuitive Fine-tuning (IFT) to integrate SFT and RLHF into a single process. IFT captures LMs’ intuitive sense of the entire answers through a temporal residual connection, but it solely relies on a single policy and the same volume of non-preference-labeled data as SFT. Our experiments show that IFT performs comparably or even superiorly to sequential recipes of SFT and some typical alignment methods across several tasks, particularly those requires generation, reasoning, and fact-following abilities. An explainable Frozen Lake game further validates the effectiveness of IFT.
Intuitive Fine-Tuning:
Towards Unifying SFT and RLHF into a Single Process
Ermo Hua1, Biqing Qi1,2, Kaiyan Zhang1,2, Yue Yu1 Ning Ding1,2, Xingtai Lv1,2, Kai Tian1,2, Bowen Zhou1,∗ 1Department of Electronic Engineering, Tsinghua University, Beijing, China 2Frontis.AI, Beijing, China hem23@mails.tsinghua.edu.cn, zhoubowen@tsinghua.edu.cn
1 Introduction
Large Language Models (LLMs) have demonstrated remarkable powerful potential across various downstream tasks after pre-training on large-scale corpora. However, their instruction-following skills and trustworthiness still fall short of expectations. Therefore, algorithms such as Supervised Fine-Tuning (SFT) and Reinforcement Learning from Human Feedback (RLHF) (Ziegler et al., 2019; Ouyang et al., 2022; Lee et al., 2023) are utilized to further enhance LLMs’ abilities and align them better with human preferences.
Considering the limited effectiveness of SFT and the high cost of data construction and training computation for RLHF, these two methods are often combined to leverage their respective strengths. Unfortunately, they can only be implemented as a sequential recipe constrained by the paradigm gap between SFT and early RLHF methods, stemming from differences in loss functions, data formats, and the requirement for auxiliary models.
Figure 1: Comparison of Alignment Methods. IFT conducts alignment solely relying on positive samples and a single policy, starting from a pre-trained base model. IFT not only enjoys the second highest efficiency only lower than SFT, but also showcases comparable or even superior performance to other alignment algorithms.
Recently, Direct Preference Optimization (DPO) (Rafailov et al., 2024) was proposed to integrate Reward Modeling and Policy Optimization into one single procedure using a loss function derived from Proximal Policy Optimization (PPO) (Schulman et al., 2017), which demonstrates the potential to unify SFT and RLHF for the first time.
Henceforth, many extended methods have been tried to realize this objective by bridging the gap between SFT and DPO. Some of them (Ethayarajh et al., 2024; Hong et al., 2024) aim to transform the contrastive loss of DPO into a SFT-like cross-entropy loss, learning positive samples similar to SFT while unlearning negative samples resort to Unlikelihood Training (Welleck et al., 2019). Some others get rid of the preference-labeling process before training, switching to collect samples and labels/rewards in an online manner (Liu et al., 2023; Yuan et al., 2024; Guo et al., 2024; Calandriello et al., 2024), or just treating the SFT targets and online policy generations as positive and negative samples respectively (Chen et al., 2024; Mitra et al., 2024; Liu et al., 2024). Nevertheless, preference-labeled pairwise data is still essential in these methods, and the need for reference model only becomes unnecessary in some cases. Thus the core differences between SFT and RLHF are not eliminated thoroughly. To address this challenging issue, a deeper and more unified understanding of SFT and RLHF are needed.
In this paper, we attempt to explain the similarities and differences between SFT and RLHF by defining Preference Estimation and Transition Optimization in terms of state-action pairs within the Markov Decision Process (MDP) framework. Through this modeling, we demonstrate that SFT is simply a specialized case with inferior estimation and optimization among all RLHF methods. To be specific, classical RLHF methods usually sample the entire answer of policy based solely on the initial instruction, and align the preference reflected by the sampled sentence with human. In contrast, SFT only samples a single token based on intermediate state of target answer, which leads to a biased estimation of policy preference and an inferior alignment shortcut. Or put it in a metaphorical view, when preparing for an exam, it is obviously more beneficial to review reference answers after completing entire questions, rather than trying to infer each subsequent step based on the reference answer ahead and checking it immediately. It is the similar case in LMs’ training so that RLHF is more beneficial than SFT for enhancing LMs.
Depending on this understanding, we introduce a unified alignment algorithm named Intuitive Fine-Tuning (IFT), which builds policy preference estimation in an intuitive manner like human, who can have a vague sense of the complete answer after hearing a question. Through a closer estimation to truly policy preference than SFT, IFT achieves a comparable or even superior alignment performance compared to the sequential recipe of SFT and RLHF. Additionally, IFT requires only a single policy model, and the same volume and format of data as SFT, enjoying both data and computation efficiency. These characteristics also enable IFT to be implemented in domains where preference data is unavailable or expensive to collect.
Our main contribution are three folds: Firstly, by modeling within the MDP framework, we explain the similarities and differences between SFT and two fundamental RLHF methods, PPO and DPO. Secondly, building on this modeling, we introduce Intuitive Fine-tuning (IFT), a deeply unified version of SFT and RLHF, which enjoys similar efficiency to SFT, while having a closer performance to RLHF. Finally, through experiments in several benchmarks, we validate that IFT can perform comparably or even superiorly to different recipes of SFT and existing RLHF methods. In particular, IFT demonstrates significantly stronger reasoning and fact-following abilities than chosen baselines. An explainable toy-setting named Frozen Lake is also used for demonstrating the effectiveness of IFT.
2 Related Work
Classical Reinforcement learning (RL) has been demonstrated strong performance in various sequential decision-making and optimal control domains, including robotics (Levine et al., 2018), computer games (Vinyals et al., 2019) and others (Guan et al., 2021). There are two main categories of RL algorithms: value-based and policy-based, depending on whether they learn a parameterized policy. Value-based RL aims to fit an value function defined by Bellman Equation, containing methods such as Monte-Carlo (MC) Learning (Lazaric et al., 2007) and Temporal Difference Learning Sutton (1988); Seijen and Sutton (2014). However, value-based methods struggle in continuous or large discrete space for its greedy objective. Thus, policy-based methods were introduced to model the decision-making process using a parameterized policy. As one of its best-known algorithms, Proximal Policy Optimization (PPO) (Schulman et al., 2017) is widely used in various domains, including Natural Language Processing (NLP).
Alignment for LMs has emerged as a crucial task these years, which adjusts the LMs’ generation distribution in line with human preferences. Building on the Bradley-Terry (BT) (Bradley and Terry, 1952) Model, current alignment methods typically train models to distinguish positive and negative answers given the same instruction using RLHF/RLAIF (Ziegler et al., 2019; Ouyang et al., 2022; Lee et al., 2023). While PPO remains the primary algorithm for alignment, its high demands for computation and memory hinders its broader use. As a result, DPO (Rafailov et al., 2024) proposes training a model to serve both as a policy model and a reward model, unifying reward modeling and policy training into a single process. Without sacrificing performance, DPO decrease the costly consumption of PPO through directly value iteration similar to a preference-based format of MC instead of TD. Nevertheless, the expensive preference-labeling process and the memory-hogging pair-wise data is still required.
Improved Versions of DPO come out one after another. Efforts such as (Liu et al., 2023; Yuan et al., 2024; Guo et al., 2024; Calandriello et al., 2024; Chen et al., 2024; Mitra et al., 2024; Liu et al., 2024) focus on online sampling and automated label/reward collection, reducing the manual cost required for alignment. Methods like (Ethayarajh et al., 2024; Hong et al., 2024) aim to reduce DPO’s dependency on SFT warm-up by transforming its loss functions and data format into a SFT manner. These algorithms handle positive and negative samples using SFT objective and Unlikelihood Training (Welleck et al., 2019), respectively. However, the actual volume of training data is not decreased in these methods. Also, GPU-memory-consuming pair-wise data is still required, while the need for a reference model and preference-labeling for the entire answer trajectory is only eliminated in limited cases.
3 Preliminaries
3.1 MDP in Language Models
The MDP applied to LMs can be formally described as a tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{T},r,\rho_{0})$, where $\mathcal{S}$ is the state space comprising ordered permutations of vocabularies, $\mathcal{A}$ is the action space consisting of vocabularies defined by the tokenizer, $\mathcal{T}$ is the transition matrix indicating token generation probabilities for a given state, $r$ represents rewards for state-action pairs, and $\rho_{0}$ is the initial state typically based on given instructions. See more details in Appendix A.1.
The primary objective of Language Modeling is to train a policy $\pi_{\theta}$ with $\mathcal{T}_{\theta}$ to mimic a human policy $\pi^{*}$ with $\mathcal{T}^{*}$, aiming for the two transition matrix to become identical:
$\forall s\in\mathcal{S},a\in\mathcal{A}:\mathcal{T}_{\theta}(a|s)\rightarrow\mathcal{T}^{*}(a|s)$ (1)
This process can also be expressed using another state-state transition matrix $T$:
$\forall~{}s,s^{\prime}\in S:T_{\theta}(s^{\prime}|s)\rightarrow T^{*}(s^{\prime}|s)$ (2)
where $T$ is equivalence to $\mathcal{T}$, but instead, indicating the transition probability from one state to another.
3.2 Preference Estimation
We define the preference $\mathcal{P}$ of policy $\pi$ given an initial instruction $\rho_{0}$ as a mapping:
$\mathcal{P}(\rho_{0}):\rho_{0}\rightarrow[\pi(\rho_{0}),\pi(s_{1}),\pi(s_{2}),...]$ (3)
where $s_{i+1}=[s_{i},a_{i}]$ , $a_{i}=\pi(s_{i})$ and $s_{0}=\rho_{0}$.
During alignment, the model preference gradually approaches the human preference:
$\mathcal{P_{\theta}}(\rho_{0})\rightarrow\mathcal{P^{*}}(\rho_{0})$ (4)
$\mathcal{P_{\theta}}(\rho_{0}):\rho_{0}\rightarrow[\pi_{\theta}(\rho_{0}),\pi_{\theta}(s_{1}^{\theta}),\pi_{\theta}(s_{2}^{\theta}),\ldots]$ (5)
$\mathcal{P^{*}}(\rho_{0}):\rho_{0}\rightarrow[\pi^{*}(\rho_{0}),\pi^{*}(s_{1}^{*}),\pi^{*}(s_{2}^{*}),\ldots]$ (6)
As the truly preferences are difficult to obtain, alignment is usually conducted based on the Preference Estimation of model and human, denoted as $\hat{\mathcal{P}_{\theta}}$ and $\mathcal{\hat{P^{*}}}$ respectively. The estimations from some common methods are listed in table 1.
To make preference optimizable, every policy’s preference can also be expressed as follows:
$\mathcal{P}(\rho_{0})=\{\mathcal{T}(a|s)|\forall a\in\mathcal{A},s\in\mathcal{S}_{\rho_{0}}\}$ (7)
Here, $\mathcal{S}_{\rho_{0}}$ denotes a conditional state space that constrained by the initial state $\rho_{0}$, within which each state can only be initially derived from $\rho_{0}$. Consequently, the model preference can be optimized through transition matrix, named Transition Optimization by us.
Figure 2: The Training Paradigm of Different Methods. Symbol * and $\theta$ denote human and model respectively, with $a_{i}^{*}=\pi^{*}(s_{i}^{*})$ and $s_{i+1}^{*}=[s_{i}^{*},a_{i}^{*}]$, similarly for $\theta$. SFT uses priors deviating from model distribution, resulting in a more biased estimation of model preferences compared to PPO and DPO. IFT achieves a closer estimation than SFT by utilizing temporary residual connections, while maintaining the data and computational efficiency as SFT.
3.3 Transition Optimization
Ideally, we want to align the state-action transition matrix between model and human in a $\rho_{0}$-constrained state space:
$\forall a\in\mathcal{A},s\in\mathcal{S}_{\rho_{0}}:\mathcal{T}_{\theta}(a,s)\rightarrow\mathcal{T^{*}}(a,s)$ (8)
which is equivalent to the following format expressed by state-state transition matrix:
$\forall s\in\mathcal{S}_{\rho_{0}}:T_{\theta}(s,\rho_{0})\rightarrow T^{*}(s,\rho_{0})$ (9)
However, considering the limited data, only matrix elements representing state-action/state-state pairs contained in the dataset would be aligned. Given a target sample with instruction $\rho_{0}$ and length-$N$ answer, the objective would be:
$\forall a\in\mathcal{A},n\in[0,N]:\mathcal{T}_{\theta}(a,s_{n}^{*})\rightarrow\mathcal{T^{*}}(a,s_{n}^{*})$ (10)
Or equivalent to:
$\forall n\in[0,N]:T_{\theta}(s_{n}^{*},\rho_{0})\rightarrow T^{*}(s_{n}^{*},\rho_{0})$ (11)
$T_{\theta}(s_{n}^{*}|\rho_{0})=\prod\limits_{i=0}^{n}\mathcal{T}_{\theta}(a_{i}^{*}|s_{i}^{*})$ (12)
$T^{*}(s_{n}^{*}|\rho_{0})=\prod\limits_{i=0}^{n}\mathcal{T^{*}}(a_{i}^{*}|s_{i}^{*})$ (13)
where $s_{0}^{*}=\rho_{0}$ and $T^{*}(\rho_{0}|\rho_{0})=T_{\theta}(\rho_{0}|\rho_{0})=1$.
Consequently, the loss function can be derived from the disparities of the transition matrices between model and human.
3.4 From SFT to RLHF
We reformulate SFT, PPO and DPO using the aforementioned framework, detailed in table 1. A more comprehensible version is presented in figure 2. To compare the differences between them, we begin by introducing a fundamental theorem and corollary:
Method Preference Estimation Transition Optimization $\hat{s_{n}^{*}}$ in $\hat{\mathcal{P^{*}}}$ $\hat{s_{n}^{\theta}}$ in $\hat{\mathcal{P_{\theta}}}$ Truly $s_{n}^{*}$ $s_{n}^{\theta}$ $T_{\theta}(s_{n}^{*},\rho_{0})\rightarrow T^{*}(s_{n}^{*},\rho_{0})$ SFT $s_{n}^{*}$ $s_{n}^{*}$ $T_{\theta}(s_{n}^{*},s_{n-1}^{*})\rightarrow T^{*}(s_{n}^{*},s_{n-1}^{*})$ PPO $s_{n}^{\theta}$ $s_{n}^{\theta}$ $T_{\theta}(\hat{s_{n}^{*}},s_{n-1}^{\theta})\rightarrow T^{*}(\hat{s_{n}^{*}},s_{n-1}^{\theta})$ DPO online $s_{n}^{*}$ $s_{n}^{\theta}$ $T_{\theta}(s_{n}^{*},\rho_{0})\rightarrow T^{*}(s_{n}^{*},\rho_{0})$ offline $s_{n}^{*}$ $s_{n}^{\theta^{-}}$ $\hat{T_{\theta}}(s_{n}^{*},\rho_{0})\rightarrow T^{*}(s_{n}^{*},\rho_{0})$
Table 1: Reformulation of SFT, PPO and DPO
Theorem Given a set of events $\mathcal{Z}$, the probability of any event $z\in\mathcal{Z}$ is between 0 and 1:
$\forall z\in\mathcal{Z}:0\leq P(z)\leq 1$ (14)
If all events are mutually independent, the sum of their probabilities equals 1:
$1=\sum_{z\in\mathcal{Z}}P(z)$ (15)
The event $z^{*}$ with the highest probability has a probability greater than or equal to any other event:
$\forall z\in\mathcal{Z}:0\leq P(z)\leq P(z^{*})\leq 1$ (16)
Corollary LMs consistently assign higher probabilities to their own greedy predictions than to human preference:
$\forall s\in S:\mathcal{T}_{\theta}(\pi^{*}(s),s)\leq\mathcal{T}_{\theta}(\pi_{\theta}(s),s)\leq 1$ (17)
thus LMs tend to assign higher probabilities to its own generation than to target answer given the same initial instruction:
$\forall n\in N:T_{\theta}(s_{n}^{*},\rho_{0})\leq T_{\theta}(s_{n}^{\theta},\rho_{0})\leq 1$ (18)
where $N$ represents the length when the generation reaches the EOS token or the truncation length.
SFT provides an unbiased estimation of human preference, but a biased estimation for model:
$\hat{\mathcal{P_{\theta}}}(\rho_{0}):\rho_{0}\rightarrow[\pi_{\theta}(\rho_{0}),\pi_{\theta}(s_{1}^{*}),\pi_{\theta}(s_{2}^{*}),\ldots]$ (19)
which is caused by wrong prior state when predicting each subsequent token. Consequently, the Transition Optimization objective of SFT:
$T_{\theta}(s_{n}^{*},s_{n-1}^{*})\rightarrow T^{*}(s_{n}^{*},s_{n-1}^{*})$ (20)
secretly sets $T_{\theta}(s_{n-1}^{*},\rho_{0})=1$ during aligning $T_{\theta}(s_{n}^{*},\rho_{0})$ with $T^{*}(s_{n}^{*},\rho_{0})$. This makes an overestimation of the transition probabilities and preference of model, leading to an inferior optimization progress in SFT. Thus RLHF is needed for further preference alignment.
PPO enjoys an unbiased estimation of model preference, while employing a progressively unbiased estimation of human preference:
$\hat{\mathcal{P^{*}}}(\rho_{0}):\rho_{0}\rightarrow[\pi^{*}(\rho_{0}),\pi^{*}(s_{1}^{\theta}),\pi^{*}(s_{2}^{\theta}),\ldots]$ (21)
Initially biased, this estimation gradually becomes unbiased as the model aligns with human preference over time.
As a reusult, PPO provides a closer approximation than SFT to the actual circumstances of model in Transition Optimization:
$T_{\theta}(\hat{s_{n}^{*}},s_{n-1})\rightarrow T^{*}(\hat{s_{n}^{*}},s_{n-1})$ (22)
which sets $T_{\theta}(s_{n}^{\theta},\rho_{0})=1$ and $\hat{s_{n}^{*}}=\pi^{*}(s_{n-1}^{\theta})$, at the expense of preference-labeling, reward modeling and online sampling for estimating $\pi^{*}(s_{n-1}^{\theta})$.
DPO theoretically achieves the best estimation across all scenarios, even without reward modeling. However, obtaining pairwise preference data online is costly, as it requires real-time negative sampling from model and preference labeling by human. Thus, mainstream implementations often rely on off-policy negative samples out-of-distribution from the optimized model, which may yield unstable and sub-optimal results due to biased preference estimation and inferior transition optimization.
4 Method
While SFT is data and computation-efficient, it has an inferior approximation for both Preference Estimation and Transition Optimization. In the other side, RLHF (represented by PPO and DPO) enjoys better approximation at the expense of preference data construction. We hope to make good use of their strength, using solely the target data as SFT but having similar approximation as RLHF.
4.1 Intuitive Preference Estimation
A key distinction between SFT and RLHF lies in whether the full distribution of model preference for each initial instruction is sampled. Contrasted to RLHF, the intermediate state of the target answer used for prior in SFT may be far away from the model preference, leading to inferior outcomes.
To obtain a state estimation $\hat{s_{i}^{\theta}}$ closer to model preference, we introduce a model-based distribution disturbance function $\delta_{\theta}$ for the biased state:
$\hat{s_{i}^{\theta}}=\delta_{\theta}(s_{i}^{*})\\ =(1-\lambda)s_{i}^{*}+\lambda\pi_{\theta}(s_{i-1}^{*})$ (23)
which can also be interpreted as a temporal residual connection. Through this approach, model can predict not only the next token from intermediate state of target answer, but also develop an intuitive sense to the entire answer generation solely based on the initial instruction, deriving more accurate Preference Estimation for model:
$\hat{\mathcal{P_{\theta}}}(\rho_{0})=[(1-\lambda)\mathcal{P_{\theta}}^{sft}+\lambda\mathcal{P_{\theta}}^{truly}](\rho_{0})$ (24)
4.2 Dynamic Relation Propagation
With improved Preference Estimation, we achieve a Transition Optimization process closer to the original objective:
$\forall n\in[0,N]:\hat{T_{\theta}}(s_{n}^{*},\rho_{0})\rightarrow T^{*}(s_{n}^{*},\rho_{0})$ (25)
$\hat{T_{\theta}}(s_{n}^{*},\rho_{0})$ $=\prod\limits_{i=0}^{n-1}T_{\theta}(s_{i+1}^{*},\hat{s_{i}^{\theta}})$ (26)
which can be optimized by the following loss function that quantifies the disparities of transition between model and human:
$\mathcal{L}(\hat{T_{\theta}}(s_{n}^{*},\rho_{0}))=\log\frac{\hat{T_{\theta}}(s_{n}^{*},\rho_{0})}{T^{*}(s_{n}^{*},\rho_{0})}$ (27)
We make the same hypothesis as SFT that the optimization objective of each target intermediate state has a probability equal to 1, so that $\forall n\in[0,N]$:
$T^{*}(s_{n}^{*},\rho_{0})=1=T^{*}(s_{N}^{*},\rho_{0})$ (28)
Then, the loss function can be reformulated as:
$\mathcal{L}(\hat{T_{\theta}}(s_{n}^{*},\rho_{0}))=-\sum\limits_{i=n}^{N}\log\mathcal{T}_{\theta}(a_{i}^{*},\delta_{\theta}(s_{i}^{*}))$ (29)
which becomes easier to be implemented in parallel. See Appendix A.2 for complete derivation.
At the same time, we also prove that the objective optimized by this loss function implicitly satisfies the Bellman Equation in the on-policy case:
$V_{\theta}(\hat{s_{n}^{\theta}})=\exp\bigg{(}-\mathcal{L}\big{(}\hat{T_{\theta}}(s_{n}^{*},\rho_{0})\big{)}\bigg{)}$ (30)
The derivation is in Appendix A.3. This proof guarantees the optimization process closer to RLHF. Additionally, it ensures the optimized objective not only reflects the prediction accuracy of current token, but also considers the influence of the current selection on future generation, helping model acquire an intuitive understanding of generation, and a better causality for reasoning and fact-following. Additionally, a decay factor can be incorporated, as in the typical Bellman Equation, to ensure effectiveness in long trajectories.
5 Experiments
We conduct experiments mainly on NLP tasks. Considering the absence of an optimal policy in NLP, we also utilize the Frozen Lake environment for further validation.
5.1 Settings
Datasets. We select UltraChat-200k (Ding et al., 2023) and UltraFeedback-60k (Cui et al., 2023) as single-target and pair-wise dataset respectively.
Models. We conduct experiments on Mistral-7B-v0.1 (Jiang et al., 2023) and its fine-tuned version Mistral-7B-sft-beta (Tunstall et al., 2023). The former serves as the base model, while the latter has been fine-tuned on the UltraChat-200k.
Scenarios. We consider two different training scenarios, one is in RLHF-only mode and the other uses sequential recipe of SFT and RLHF. In the first scenario, alignment is conducted directly from base model Mistral-7B-v0.1 using UltraFeedback. In order to ensure balanced data volume between different method, we randomly sample 60k data from UltraChat as supplementary for SFT and IFT, for only the target data are utilized in these two methods. The second scenario is commonly seen, where SFT and RLHF is employed sequentially. For this scenario, we use Mistral-7B-sft-beta as start-point, which has been fine-tuned with UltraChat using SFT. Then we fine-tune it further with UltraFeedback using RLHF.
Baselines. Our main baselines are SFT and DPO (Rafailov et al., 2024), with PPO excluded due to computational limitations. Additionally, we incorporate an improved versions of DPO, named ORPO (Hong et al., 2024), who claims to achieve alignment directly without SFT and reference model, aligning well with our goals. In addition to reproducing the algorithms mentioned above, we also consider Zephyr-7B-beta (Tunstall et al., 2023) and Mistral-ORPO-alpha (Hong et al., 2024), two open-source checkpoints that utilize sequential and direct recipes respectively. Both of them used start-point models and datasets similar to ours.
Benchmarks. We consider two types of benchmarks. One is from the widely used Open-LLM LeaderBoard, which contains ARC-Challenge (25-shot) (Clark et al., 2018), MMLU (5-shot) (Chung et al., 2024), TruthfulQA (0-shot) (Lin et al., 2021), WinoGrande (5-shot) (Sakaguchi et al., 2021), and GSM8K (5-shot) (Cobbe et al., 2021). The other is LM-based evaluation, including Alpaca-Eval and Alpaca-Eval-2 (Dubois et al., 2024). We utilize chat template for all benchmarks to obtain a more accurate evaluation for chat models.
5.2 Main Results
Effectiveness on Sequential Recipe. In this scenario, IFT demonstrates good performance across benchmarks having standard answers or not, see more details in table 2 and 4. On Open-LLM Leaderboard, IFT showcases the best average capabilities across all tasks, excelling particularly in tasks requiring generation, reasoning and fact-following abilities, such as TruthfulQA and GSM8K. However, IFT has a relatively large gap between DPO in multi-choice tasks like ARC-Challenge and MMLU. We will explained the reasons later. When evaluated for instruction-following and question-answering abilities as judged by GPT-4, IFT’s performance is comparable to that of the chosen baselines. Remarkably, IFT achieves these results using the least amount of data and computational resources among all the methods tested.
Method Alpaca-Eval Alpaca-Eval-2 win-rate lc win-rate win-rate lc win-rate Mistral-7B 24.72 11.57 1.25 0.35      with UltraFeedback-60k + SFT 82.56 78.32 7.09 8.67 + DPO 74.00 73.12 9.73 8.58 + ORPO 85.14 76.60 8.82 12.34 + IFT 85.18 78.78 9.95 13.27 Mistral-ORPO-$\alpha$ 87.92 11.33      with UltraChat-200k + UltraFeedback-60k sequentially + SFT 86.69 77.96 4.08 6.43 + SFT + SFT 86.34 76.98 4.55 7.14 + SFT + DPO 91.62 81.54 10.08 13.72 + SFT + ORPO 86.26 79.67 7.40 12.27 + SFT + IFT 88.37 81.29 10.26 14.34 Zephyr-7B-$\beta$ 90.60 10.99
Table 2: Evaluation on LLM-based Benchmarks. In the typical sequential setting, SFT+DPO still achieves the highest scores, although IFT performs comparably. However, when aligning directly from the base model, IFT performs not only superiorly to other methods but also on par with sequential recipes. As for open-source models, we directly present its publicly available score.
Method Data Volume GPU Memory Training Duration SFT v m t DPO 2v $\sim$2m $\sim$3t ORPO 2v $\sim$1.5m $\sim$2t IFT v m $\sim$1.5t
Table 3: Cost Estimation for Different Methods. We use the cost of SFT as the benchmark, where v, m, and t do not represent exact quantity units. We assume backward propagation needs higher cost than forward propagation in terms of GPU Memory and Training Duration.
Method ARC ARC-Gen MMLU TruthfulQA WinoGrande GSM8K Avg. Mistral-7B 53.07 73.04 59.14 45.29 77.58 38.89 54.79      fine-tuning with UltraFeedback-60k + SFT 56.49 74.00 60.44 55.57 77.90 42.84 58.65 + DPO 61.86 73.54 61.02 47.98 76.64 43.89 58.28 + ORPO 56.66 73.98 60.57 51.77 77.19 42.30 57.70 + IFT 56.74 74.15 60.49 57.65 78.45 44.73 59.61 Mistral-ORPO-$\alpha$ 57.25 73.72 58.74 60.59 73.72 46.78 59.41      fine-tuning with Ultrachat-200k + UltraFeedback-60k sequentially + SFT 57.68 72.87 58.25 45.78 77.19 40.94 55.97 + SFT + SFT 58.10 72.61 58.40 48.59 76.80 43.06 56.99 + SFT + DPO 63.91 73.98 59.75 46.39 76.06 41.47 57.52 + SFT + ORPO 58.45 73.21 58.80 50.31 76.45 42.76 57.35 + SFT + IFT 58.36 73.38 58.45 52.39 78.06 43.82 58.22 Zephyr-7B-$\beta$ 67.41 72.61 58.74 53.37 74.11 33.89 57.50
Table 4: Evaluation on Open-LLM Leaderboard with chat template. When fine-tuning with the same recipe, IFT achieves the highest average score across all methods. Directly conducting alignment using IFT showcases the best performance in all recipes with the least data and computation. In particular, IFT performs better in tasks requiring generation, reasoning and fact-following abilities, but falling behind in multi-choice tasks like ARC and MMLU.
Effectiveness on RLHF-only Recipe. IFT not only maintains the performance advantages compared with other baselines in this setting, as seen in the sequential scenario. But also, IFT performs comparably or even superiorly to many method in sequential recipe. While DPO tends to fail under this setting, ORPO remains competitive in its open-source model. However, when constrained in the same experiment setting, the performance of ORPO becomes worse than IFT. Additionally, the reliance on preference data makes ORPO more costly in terms of negative sampling, preference labeling, and GPU memory consumption. Consequently, IFT stands out as a more efficient and cost-effective alternative in this context.
Multi-Choice vs. Generation. When evaluated on Open-LLM Benchmark, IFT performs better on generation tasks but struggles with multi-choice tasks, whereas DPO exhibits the opposite performance. This discrepancy may be partially due to the different evaluation metrics between these two type of tasks, and different training objective between IFT and DPO. Multi-choice tasks are evaluated using the log-likelihood of model for each entire golden answer, while generation tasks require the model to build the final answer based on the selected token by itself ahead. So generation better reflects the causality and reasoning abilities than multi-choice. While DPO only focus on aligning the mapping between instructions and entire answers, IFT not only addresses this issue, but also places greater emphasis on the causal relationship between tokens. As a result, DPO tends to be expert on distribution mapping tasks like multi-choice, but IFT tends to perform better on tasks that requires model to explore token by token. Considering this, we changed ARC-Challenge into a generation task, during which the question and candidate answers would be displayed beforehand and answer would be extract from model’s generation. Without changing the distribution of the benchmark, IFT showcases advantage compared with any other method in this setting. Overall, IFT maintains the best balance across different tasks, achieving the highest average score.
Objective Trade-off between SFT and RLHF. Traditional RLHF methods deliver excellent alignment performance, particularly in enhancing the instruction-following ability of language models, as showed in table 2. However, fitting the different objectives of SFT and RLHF involves many hyper-parameter trade-offs. This phenomenon is also observed in table 4, where the models trained by sequential recipe of SFT and other RLHF methods showcase obvious inferior results on Open-LLM Leaderboard even worse than SFT alone. Avoiding this trade-off, ORPO and IFT can achieve better performance by directly conducting alignment on the base model.
Efficiency and Scaling Potential of IFT. Although IFT achieves comparable or superior performance to other methods, it also boasts high efficiency in many aspects. Like SFT and ORPO, IFT does not require a reference model, which conserves GPU memory and computational resources. Most importantly, IFT and SFT are the only methods that conduct alignment without preference data, offering significant benefits as follows. Firstly, this characteristic eliminates the need for synchronous storage and computation of pairwise data on the GPU, thereby reducing memory consumption and training duration. Secondly, negative sampling from models and human preference-labeling are no longer necessary, eliminating the highest cost associated with alignment, which has been a discarded but fundamental challenge in research so far. Furthermore, using only the target answer brings the potential for scaling in alignment process, mirroring the core benefits found in pre-training.
5.3 Analysis in Frozen-Lake Environment
Figure 3: The Frozen Lake Game. IFT performs much better than SFT and ORPO, slightly worse than DPO.
As scores on Open-LLM Leaderboard only partially reflect models’ performance, and GPT-4 inadequately models human language generation, further comparison to a truly optimal policy is necessary. Given the difficulty of obtaining an optimal policy representing human language, we validate our algorithm in a simplified setting called Frozen Lake (Farama, 2023). In this environment, an agent attempts to find a gift on a nearly frozen lake with several holes, terminating the game upon finding the gift or falling into a hole. The limited number of states and actions allows the optimal policy to be easily derived using classical RL methods.
To simulate parameterized policy alignment, we employ a two-layer fully connected neural network and design the environment with one optimal and one sub-optimal trajectory. The optimal parameterized policy is trained using the previously obtained optimal state-action probabilities, and various fine-tuning methods from LMs are compared. We evaluate performance by measuring the MSE distance between the optimal and trained policy parameters.
IFT achieves a significantly better policy than SFT and ORPO, although it performs slightly worse than DPO. This is partly because, in terms of comparing how closely the explored grid aligns with the agent’s preference, the order is DPO > IFT > ORPO > SFT. Although ORPO also considers the negative trajectories sampled from policy, its direct incorporation of SFT loss with a fusion coefficient deviates its preference estimation, partially diminishing its effectiveness. Additionally, DPO, ORPO and IFT explore more grids than SFT, which also helps the agent develop a better understanding of the environment.
6 Conclusion
In this paper, we first interpret SFT and some typical RLHF methods into a unified framework using Preference Estimation and Transition Optimization. We then introduce an efficient and effective method called Intuitive Fine-Tuning (IFT), which achieves alignment directly from the base model using non-preference-labeled data. Finally, experiments on widely used benchmarks in NLP setting and in Frozen Lake environment demonstrate the competitive performance of IFT.
Limitations and Future Work. Our validation of IFT is limited to the fine-tuning setting, where data volume is constrained, leaving the scalability of IFT unexplored. Additionally, we primarily use Mistral-7B for baseline testing, and the generalization of IFT to larger and more diverse models requires exploration in the future.
Acknowledgments
Special thanks to Yihao Liu, Che Jiang, and Xuekai Zhu for contributing to idea formulation, Jingkun Yang, and Xuanqi Dong for their assistance with mathematical definitions, and Hong Liu, and Chushu Zhou for insightful discussions during the experiments.
Appendix A Appendix
A.1 MDP In LMs
$\mathcal{M}=(S,A,\mathcal{T},r,\rho_{0})$:
  • $A$, the concrete action space, consisting of $N_{A}$ vocabularies as defined by the tokenizer.
  • $S$, the concrete state space, comprising $N_{S}=(N_{A})^{N}$ elements related to sequence length $N$. Each state represents a ordered permutation of vocabularies.
  • $\rho_{0}$, the initial state of each generation, typically refers to the given instruction;
  • $\mathcal{T}\in R^{N_{S}\times N_{A}}$, the state-action transition matrix of a given policy, indicating the probability of generating each token given different states;
  • $r$, the reward assigned to a particular state-action pair.
A.2 Loss Function of IFT
$\mathcal{L}(\hat{T_{\theta}}(s_{n}^{*},\rho_{0}))$ $=\log\frac{\hat{T_{\theta}}(s_{n}^{*},\rho_{0})}{T^{*}(s_{N}^{*},\rho_{0})}$ (31)
$=\log\frac{1}{\hat{T_{\theta}}(s_{N}^{*},s_{n}^{*})}$
$=-\log\prod\limits_{i=n}^{N}\mathcal{T}_{\theta}(a_{i}^{*},\delta_{\theta}(s_{i}^{*}))$
$=-\sum\limits_{i=n}^{N}\log\mathcal{T}_{\theta}(a_{i}^{*},\delta_{\theta}(s_{i}^{*}))$
A.3 Proof for Bellman Equation
$\exp\big{(}-\mathcal{L}(\hat{T_{\theta}}(s_{n}^{*},\rho_{0}))\big{)}$ (32)
$=\mathcal{T}_{\theta}(a_{n}^{*},\delta_{\theta}(s_{n}^{*}))\bigg{(}\sum\limits_{n+1}^{N}\mathcal{T}_{\theta}(a_{i}^{*},\delta_{\theta}(s_{i}^{*}))\bigg{)}$
$=\max\limits_{a}\bigg{[}\mathcal{T}_{\theta}(a,s_{n}^{*})\big{(}r+\gamma V(\hat{s^{\theta}_{n+1}})\big{)}\bigg{]}$
$=V_{\theta}(\hat{s_{n}^{\theta}})$
, where $r=(1-\gamma)V(\hat{s^{\theta}_{n+1}})$.